# A geometric approach to informed MCMC sampling

### v1

@@ 7 Examples
Sampler $E(X_{1})$ $E(X_{2})$ MSJD mESS ESS $(X_{1})$ ESS $(X_{2})$
RW -0.004 -0.004 0.916 11911 12224 11437
GMC1(RW) - 6.729 - 6.721 29.586 17520 15132 15549
GMC2(RW) - 5.025 - 5.035 33.352 23740 20371 18756
GMC3(RW) - 4.675 - 4.705 0.957 1152 16578 17133
@@
We analyze the Pima Indian data set by fitting a Bayesian logistic regression model using the RWM, independent MH algorithms, the MALA, the MMALA, and their geometric MH variants. We take $\mu_{0}=0_{8}$, a vector of zeros, and $\Sigma_{0}=10^{3}I_{8}$ for the prior density of $\beta$. We ran the algorithms for 100,000 iterations all started at $\mu_{0}$. Let $\hat{\beta}$ be the MLE of $\beta$. We obtain $\hat{\beta}$ by fitting the glm function of R with the logit link. Let $\hat{\Sigma}=(-\nabla^{2}\log\psi(\beta))^{-1}|_{\beta=\hat{\beta}}$, the generalized observed Fisher information matrix. For the RWM, we use a normal proposal with covariance matrix $\Sigma_{f}=0.3\hat{\Sigma}$ to get an acceptance rate of around 50%. For the independent MH, we use a normal proposal with mean $\hat{\beta}$ and covariance matrix $\hat{\Sigma}$ resulting in an acceptance rate of around 80%. For constructing MALA and MMALA, we need the first and higher-order derivatives of the log target density $\psi$. Indeed, the proposal density of MALA is $\phi(\beta^{\prime};\beta+h\nabla\log\psi(\beta)/2,hI_{d})$ for some step-size $h$ and $\beta$ is the current state of the Markov chain. It turns out that $\nabla\log\psi(\beta)=W^{\top}(z-\xi)-\Sigma_{0}^{-1}(\beta-\mu_{0}),\nabla^{2}\log\psi(\beta)=-W^{\top}\Lambda W-\Sigma_{0}^{-1}$ and $\frac{\partial\nabla^{2}\log\psi(\beta)}{\partial\beta_{j}}=-W^{\top}\Gamma^{j}X$, where $W$ is the $m\times p$ matrix of covariates, $\xi=(\xi_{1},\dots,\xi_{m})$, $\Lambda$ is the $m\times m$ diagonal matrix with $i$th diagonal element $\xi_{i}(1-\xi_{i}),i=1,\dots,m$, and $\Gamma^{j}$ is the $m\times m$ diagonal matrix with $i$th diagonal element $\xi_{i}(1-\xi_{i})(1-2\xi_{i})w_{ij},i=1,\dots,m$. For MALA and MMALA we take $h=0.01$ and $h=2$, respectively resulting in around 50% acceptance rates.
@@
Table 3: First eight autocorrelations for different samplers applied to the Pima Indian data
@@
where $c_{\gamma}=\sum_{\gamma^{\prime}\in\mathcal{N}(\gamma)}\psi(\gamma|z)$. For implementing this proposed algorithm, we need to efficiently compute $\psi(\gamma|z)$ for all $\gamma^{\prime}\in\mathcal{N}(\gamma)$. Note that Zhou et al. (2022) consider only the ‘addition’ and ‘deletion’ moves in their simulation examples, greatly reducing the computational burden. Here, we use the fast Cholesky updates of Li et al. (2023) for rapidly computing $\psi(\gamma|z)$ for all $\gamma^{\prime}\in\mathcal{N}(\gamma)$. Furthermore, both Yang et al. (2016) and Zhou et al. (2022) use Zellner’s (1986) $g-$prior on $\beta$ and a different prior on $\gamma$. The $g-$prior, although a popular alternative to the independent normal prior (20b), it requires all $m\times q$ sub-matrices of $W$ have full column rank for $q\leq m-1$ and the support of the prior on $\gamma$ is restricted to models of size at most $m-1$.
@@
Our simulation experiments are conducted using 100 simulated pairs of training and testing datasets. For each simulation setting, we set $p=10000$ and $m=400$ for both training and testing data sets. Following Li et al. (2023), we choose $w$ and $\lambda$ to be $\sqrt{m}/p$ and $m/p^{2}$, respectively. The error variance $\sigma^{2}$ is set by assuming different theoretical $R^{2}$ values. While the results for $R^{2}=90\%$ are provided in Table 4, the Appendix B contains the results for $R^{2}=60\%$ and $R^{2}=75\%$. All results are based on 100 iterations of the GMC chains and 50,000 iterations of the RW chains with all chains started at the null model.
@@
We first ran the GMC1 chain with $\lambda=m/p^{2}=5.61865e-$10 and $w=\sqrt{m}/p=2.37037e-$5 (the default choices for $\lambda$ and $w$ in the function geomc.vs in the accompanying R package geommc) for $N=100$ iterations starting from the null model. Only one variable resulted in with estimated MIP above 0.5. Indeed, $\hat{\gamma}_{\text{med}}=1_{83878775}$. The $R^{2}$ value based on fitting $\hat{\gamma}_{\text{med}}$ on the response $z$ is 20.07%. Next, we ran the GMC1 chain for 100 iterations starting at the null model, but this time following Li et al. (2023), we took $\lambda=\sqrt{m}=19.20937$ (high shrinkage) and a higher value for $w=m/p=0.00046$. This time, $\hat{\gamma}_{\text{med}}$ contained four variables with $\hat{\gamma}_{\text{med}}=(1_{83878775},2_{175541357},3_{226820323},8_{27127525})$. In addition to the estimated median model $\hat{\gamma}_{\text{med}}$, we also consider a weighted average model (WAM) $\hat{\gamma}_{\text{wam}}$. For the unique MCMC samples $\{\gamma^{(n)}\}_{n=1}^{N}$ we assign the weights $\text{wt}_{n}=\psi(\gamma^{(n)}|z)/\sum_{i=1}^{N}\psi(\gamma^{(i)}|z)$ according to the marginal posterior pmf (21). Then, the approximate marginal inclusion probability for the $j$th variable is computed as $\hat{\pi}_{j}=\sum_{n=1}^{N}\text{wt}_{n}\mathbb{I}(\gamma^{(n)}_{j}=1)$ and define the WAM as the model containing variables $j$ with $\hat{\pi}_{j}>0.5.$ Based on the 100 iterations of the GMC1 chain, $\hat{\gamma}_{\text{wam}}$ consisted of five variables with $\hat{\gamma}_{\text{wam}}=(1_{83878775},2_{79769999},2_{175541357},8_{27127525},10_{10606917})$. The $R^{2}$ values for $\hat{\gamma}_{\text{med}}$ and $\hat{\gamma}_{\text{wam}}$ are 39.10% and 43.93%, respectively. We analyzed the real dataset on a Linux server equipped with 128 cores of AMD EPYC 7542 CPU and 1 TB of RAM, where it took 5.34 minutes to complete 100 iterations of GMC1 with $\lambda=\sqrt{m}$ and $w=m/p$. With these values of $(\lambda,w)$, we repeated 100 iterations of GMC1 50 times, each time with a different seed. The models $\hat{\gamma}_{\text{med}}$ and $\hat{\gamma}_{\text{wam}}$ vary in these 50 runs, suggesting that the posterior surface is highly multimodal. The range of the size of $\hat{\gamma}_{\text{med}}$ and $\hat{\gamma}_{\text{wam}}$ are (2,7) and (3, 9), respectively. Also, the ranges of $R^{2}$ values for $\hat{\gamma}_{\text{med}}$ and $\hat{\gamma}_{\text{wam}}$ are (27.29%, 46.85%) and (32.76%, 57.18%), respectively.
@@
When we ran GMC2 for 100 iterations with $\lambda=m/p^{2}$ and $w=\sqrt{m}/p$ started at the null model, both $\hat{\gamma}_{\text{med}}$ and $\hat{\gamma}_{\text{wam}}$ resulted in the empty model, although the marker with the largest MIP and the largest weighted MIP was the same SNP $1_{83878775}$ obtained by GMC1 with these choices for $(\lambda,w)$. Indeed, the largest MIP and weighted MIP were 0.47 and 0.49, respectively. Next, we ran the GMC2 chain for 100 iterations started at the null model with $\lambda=\sqrt{m}$ and $w=m/p$. Based on the 100 iterations of the GMC2 chain, $\hat{\gamma}_{\text{med}}$ consisted of eight variables with $\hat{\gamma}_{\text{med}}=(1_{83878648},2_{175541342},2_{179751269},4_{225874150},5_{21046482},\\ 8_{14610995},8_{72384070},$ $8_{115299982})$ and $\hat{\gamma}_{\text{wam}}=(1_{83878648},2_{175541342},3_{170821024},4_{237392971},8_{72384070},\\ 9_{4338284})$. The $R^{2}$ values for $\hat{\gamma}_{\text{med}}$ and $\hat{\gamma}_{\text{wam}}$ in this case are 43.09% and 49.37%, respectively. It took 11.22 minutes to complete 100 iterations of GMC2 with $\lambda=\sqrt{m}$ and $w=m/p$. Finally, based on 50 repetitions of 100 iterations of the GMC2 chain with $\lambda=\sqrt{m}$ and $w=m/p$, each time with a different seed, we observe the ranges of the size of $\hat{\gamma}_{\text{med}}$ and $\hat{\gamma}_{\text{wam}}$ are (4,17) and (2, 8), respectively. Also, the ranges of $R^{2}$ values for $\hat{\gamma}_{\text{med}}$ and $\hat{\gamma}_{\text{wam}}$ are (40.11%, 60.10%) and (28.91%, 54.61%), respectively.

### v2

@@ 7 Examples
Sampler $E(X_{1})$ $E(X_{2})$ MSJD mESS mESS/sec ESS
RW -0.004 -0.004 0.916 11911 964 (12224,11437)
GMC1(RW) - 6.729 - 6.721 29.586 17520 156 (15132,15549)
GMC2(RW) - 5.025 - 5.035 33.352 23740 17 (20371,18756)
GMC3(RW) - 4.675 - 4.705 0.957 1152 15 (16578,17133)
Example 4 (A target distribution with six modes).
The following target density (15) is from Leman et al. (2009)
$\psi(x_{1},x_{2})\propto\exp\bigg{(}\frac{-x_{1}^{2}}{2}\bigg{)}\exp\bigg{(}-\frac{((csc\leavevmode\nobreak\ x_{2})^{5}-x_{1})^{2}}{2}\bigg{)},\;-10\leq x_{1},x_{2}\leq 10.$ (15)
The marginal densities of $X_{1}$ and $X_{2}$, known up to a normalizing constant, given in Figure 6, clearly show that the target distribution has six well separated modes.
Figure 6: Desired and attained marginals from 100,000 iterations of two MCMC samplers for the six-mode target example. The first and second plots from the left show marginal histograms from the RW-within-Gibbs sampler demonstrating that the chain entirely lies in one mode. Whereas the histograms (the third and fourth plots from the left) obtained from the geometric RW-within-Gibbs sampler show that the chain successfully moves between the modes.
We first consider a RW-within-Gibbs sampler. In particular, following Leman et al. (2009), this sampler iteratively samples from $\psi(x_{1}|x_{2})$ and $\psi(x_{2}|x_{1})$ using a RWM algorithm with the normal proposal with standard deviation $0.1$. We ran this chain, denoted by RW-w-Gibbs, for 100,000 iterations starting at $(0.05,1.5)$. From Figure 6 we see that the chain is trapped in one of the modes, and has failed to converge to the target density (15).
Next, we ran another algorithm where iteratively sampling from the conditional densities $\psi(x_{1}|x_{2})$ and $\psi(x_{2}|x_{1})$ are done using the proposed geometric RW sampler. That is, in every iteration of the Gibbs sampler the two RWM steps for the conditional densities are replaced by one transition from their corresponding GMC (RW) steps with the same baseline RW densities. While the base densities are the same as the proposals in RW-w-Gibbs, a ‘non-informative’ choice, namely a normal density with mean zero and variance 900, is made for the $g$ density. This sampler, denoted by GMC(RW)-w-Gibbs, was also run for 100,000 iterations starting at $(0.05,1.5)$. From Figure 6 we see that the geometric MH-within-Gibbs algorithm has successfully accomplished the two important tasks: find all the modes, and move between them. Indeed as in Example 3, the ‘global’ moves of the geometric proposal complements the localized steps of the RW by moving between the modes.
The better mixing of the GMC(RW)-w-Gibbs chain is corroborated by much higher MSJD values given in Table 2. The geometric sampler results in estimates of the marginal means closer to the true values, whereas the estimates obtained from the RW-w-Gibbs chain are far off. As mentioned in the previous example, the ESS estimates for the RW-w-Gibbs chain can’t be trusted. Finally, we consider an algorithm where only the RW step for $\psi(x_{2}|x_{1})$ is replaced by its geometric variant mentioned above. From the histogram given in Figure S2 in the supplement, we see that this chain, denoted by GMC($X_{2}$)-w-Gibbs, also successfully moves between the modes although compared to when geometric sampler is used to sample from $\psi(x_{1}|x_{2})$, GMC($X_{2}$)-w-Gibbs results in poorer estimate of the marginal density of $X_{1}$. We also ran the Markov chain obtained by replacing only the RW step for $\psi(x_{1}|x_{2})$ with its geometric invariant. This chain remained stuck (not shown here) in a local mode for the whole 100,000 iterations.
Table 2: Results for different samplers for the six-modal target example
Sampler $E(X_{1})$ $E(X_{2})$ MSJD mESS mESS/sec ESS
RW-w-Gibbs -0.681 - 1.569 0.017 749 68 (530,1740)
GMC(RW)-w-Gibbs -0.172 --0.027 1.450 337 5 (90,973)
GMC($X_{2}$)-w-Gibbs -0.013 0.033 1.469 756 6 (373, 998)
@@
We analyze the Pima Indian data set by fitting a Bayesian logistic regression model using the RWM, independent MH algorithms, the MALA, the MMALA, and their geometric MH variants. We take $\mu_{0}=0_{8}$, a vector of zeros, and $\Sigma_{0}=10^{3}I_{8}$ for the prior density of $\beta$. We ran the algorithms for 100,000 iterations all started at $\mu_{0}$. Let $\hat{\beta}$ be the MLE of $\beta$. We obtain $\hat{\beta}$ by fitting the glm function of R with the logit link. Let $\hat{\Sigma}=(-\nabla^{2}\log\psi(\beta))^{-1}|_{\beta=\hat{\beta}}$, the generalized observed Fisher information matrix. For the RWM, we use the normal proposal with covariance matrix $\Sigma_{f}=0.3\hat{\Sigma}$ to get an acceptance rate of around 50%. For the independent MH, we use a normal proposal with mean $\hat{\beta}$ and covariance matrix $\hat{\Sigma}$ resulting in an acceptance rate of around 80%. For constructing MALA and MMALA, we need the first and higher-order derivatives of the log target density $\psi$. Indeed, the proposal density of MALA is $\phi(\beta^{\prime};\beta+h\nabla\log\psi(\beta)/2,hI_{p})$ for some step-size $h$ and $\beta$ is the current state of the Markov chain. It turns out that $\nabla\log\psi(\beta)=W^{\top}(z-\xi)-\Sigma_{0}^{-1}(\beta-\mu_{0}),\nabla^{2}\log\psi(\beta)=-W^{\top}\Lambda W-\Sigma_{0}^{-1}$ and $\frac{\partial\nabla^{2}\log\psi(\beta)}{\partial\beta_{j}}=-W^{\top}\Gamma^{j}X$, where $W$ is the $m\times p$ matrix of covariates, $\xi=(\xi_{1},\dots,\xi_{m})$, $\Lambda$ is the $m\times m$ diagonal matrix with $i$th diagonal element $\xi_{i}(1-\xi_{i}),i=1,\dots,m$, and $\Gamma^{j}$ is the $m\times m$ diagonal matrix with $i$th diagonal element $\xi_{i}(1-\xi_{i})(1-2\xi_{i})w_{ij},i=1,\dots,m$. For MALA and MMALA we take $h=0.01$ and $h=2$, respectively resulting in around 50% acceptance rates.
@@
Table 4: First eight autocorrelations for different samplers for the Pima Indian data
@@
Example 6 (Bayesian spatial GLMM).
In spatial GLMMs, conditional on the underlying latent spatial Gaussian process $\{S(l),l\in\mathcal{L}\}$, the observations $z=(z(l_{1}),\dots,z(l_{m}))$ at observed locations $(l_{1},\dots,l_{m})$ are assumed to be independent random variables following the exponential family with their (conditional) means related to $S(l)$’s via a link function. Let $S_{i}\equiv S(l_{i})$ denotes the value of the underlying process at location $l_{i},i=1,\dots,m$ and $S=(S_{1},S_{2},\dots,S_{m})^{\top}$. Denoting $z(l_{i})$ by $z_{i}$, for analyzing spatial count data, we assume that conditional on $S$, $z_{i}|s_{i}\stackrel{{\scriptstyle\mathrm{ind}}}{{\sim}}\mathrm{Poisson}(t_{i}\lambda_{i})$ where $\lambda_{i}$ is the rate parameter, the link function is $s_{i}=\log(\lambda_{i})$ and $t_{i}$ is some ‘exposure’ variable to the event, for example, the number of hours of operation, or the area for the $i$th location, $i=1,\dots,m$. Assume that $S$ has the mean $X\beta$, where $X$ is the $m\times p$ covariates matrix and $\beta\in\mathbb{R}^{p}$ is the regression parameter. Also, let the covariance matrix $\Sigma$ of $S$ be formed from an exponential correlation function that is, $\text{cov}(S(l),S(l^{\prime}))=\sigma^{2}\exp\{-\theta\|l-l^{\prime}\|+\omega\}$, where $\sigma^{2},\theta,\omega$ are the partial sill, range and relative nugget parameters, respectively, and $\|l-l^{\prime}\|$ is the Euclidean distance between $l$ and $l^{\prime}$. There are other choices of parametric correlation functions available in the literature (Diggle et al., 2003).
We consider a Bayesian analysis and assume the normal prior on $\beta$ with mean $\mu_{0}$ and covariance matrix $\sigma^{2}\Sigma_{0}$, that is, apriori $\beta|\sigma^{2}\sim N_{p}(\mu_{0},\sigma^{2}\Sigma_{0})$. The parameters $(\sigma^{2},\theta,\omega)$ are assumed apriori independent with $\sigma^{2}\sim\text{Inverse Gamma}(\alpha_{1},\gamma_{1}),\theta\sim\text{Inverse Gamma}(\alpha_{2},\gamma_{2}),$ and $\omega\sim\text{Inverse Gamma}(\alpha_{3},\gamma_{3})$ for some known hyper parameter values of $\alpha_{i}$ and $\gamma_{i}$, $i=1,2,3$. We run different MH-within-Gibbs algorithms to sample from the posterior density $\psi(S,\beta,\sigma^{2},\theta,\omega|z)$. The joint density $\psi(S,\beta,\sigma^{2},\theta,\omega|z)$ and its conditionals are derived in the supplement. From these derivations, we see that the full conditional densities of $\beta$ and $\sigma^{2}$ are normal and inverse gamma densities, respectively. Conversely, the conditional densities of $S$, $\theta$, and $\omega$ are not standard. Among these, the density of $S$ is high-dimensional, with the dimension being the same as the number of observations, $m$. We use RWM steps for sampling from the conditional densities of $\log(\theta)$ and $\log(\omega)$. For sampling from the conditional density of $S$, we consider various MH algorithms, particularly RWM, MALA, and MMALA, and their geometric variants. The supplement provides the first and higher-order derivatives of the logarithm of the conditional density of $S$ required for constructing the MALA and MMALA. Also, the order in which different variables are sampled in the Gibbs samplers are given in the supplement.
We consider the Rongelap island dataset, which consists of the measurements of $\gamma$-ray counts $z_{i}$ observed during $t_{i}$ seconds at $i$th coordinate on the Rongelap island, $i=1,\ldots,m$, with $m=157$. This data set was analyzed by Diggle et al. (1998); Evangelou and Roy (2019) among others, using a Poisson spatial model. For this dataset, $p=1$ and we assume $\mu_{0}=0$, and $\Sigma_{0}=100$. For the priors on $\sigma^{2},\theta$ and $\omega$, we assume $\alpha_{1}=\alpha_{2}=\alpha_{3}=2.04$ and $\beta_{1}=\beta_{2}=\beta_{3}=2.08$ setting their prior means and variances at 2 and 100, respectively.
For the RWM for the conditional densities of $\log(\theta)$ and $\log(\omega)$, we use normal proposals with variances 0.04 and 1, respectively to get acceptance rates of around 50%. For sampling from the conditional density of $S$ using RWM we use the normal proposal with covariance matrix $3*10^{-5}I_{157}$ and for MALA and MMALA we take $h=0.0001$ and $h=0.5$, respectively resulting in around 50% acceptance rates. We ran the algorithms for 100,000 iterations all started at $s=\log(z),\beta=\mu_{0},\sigma^{2}=\theta=\omega=2$ on a Linux server equipped with 128 AMD EPYC 7542 CPU cores and 1 TB of RAM.
For the geometric variants of the algorithms mentioned above, for sampling from the conditional density of $S$, we take $k=1$ and $g(s)=\phi_{157}(s;s_{\text{mean}},G)$, where $s_{\text{mean}}$ is the estimate of the posterior mean of $S$ obtained from the above mentioned MH-within-Gibbs sampler with the RWM step for $S$ and $G\equiv G_{\sigma^{2},\theta,\omega}=(-\nabla^{2}\log\psi(s|\beta,\sigma^{2},\theta,\omega,z))^{-1}|_{s=s_{\text{mean}}}$ evaluated at the values of $\sigma^{2},\theta,$ and $\omega$ in that iteration. Note that $G$ and hence the density $g$ change in every iteration, unlike in the previous examples.
Table 5 provides ESS, mESS, their time normalized values, and MSJD values for $S$ based on the different samplers. For ESS, we provide the minimum, median, and maximum of the 157 values corresponding to $m=157$ coordinates of $S$. Table 6 shows the first eight autocorrelations for the function $S^{\top}S$ for the different Markov chains. From Tables 5-6 we see that GMC(MMALA) has superior performance over the other algorithms in terms of autocorrelation, MSJD, median ESS and time normalized ESS, mESS values. As in Example 5, the improvement of the geometric algorithm over an uninformed base kernel like RW is much higher than that over a more informed kernel like MALA and MMALA. Also, as in Example 5, for the geometric variants of the algorithms, we used $\epsilon=0.5$, and other values remained the same as the non-geometric chains.
Table 5: Multivariate ESS and ESS (minimum, median, maximum) and their time normalized values for different samplers for the Rongelap island data
Sampler mESS ESS mESS/sec ESS/sec MSJD
RW 651 (7,110,650) 0.157 (0.002,0.027,0.157) 0.000
GMC(RW) 30077 (21535,24487,27973) 5.084 (3.640,4.139,4.729) 0.067
MALA 11573 (98,4055,49329) 2.414 (0.020,0.846,10.290) 0.010
GMC(MALA) 35927 (23312,27997,61011) 5.580 (3.621,4.349,9.476) 0.072
MMALA 9952 (6581,8207,11121) 2.021 (1.337,1.667,2.259) 0.026
GMC(MMALA) 34640 (25600,29244,34472) 7.065 (5.221,5.964,7.030) 0.075
Table 6: First eight autocorrelations for different samplers for the Rongelap island data
Sampler lag1 lag2 lag3 lag4 lag5 lag6 lag7 lag8
RW 0.998 0.996 0.994 0.992 0.990 0.989 0.987 0.985
GMC(RW) 0.586 0.349 0.212 0.129 0.088 0.061 0.038 0.023
MALA 0.926 0.862 0.803 0.751 0.704 0.659 0.617 0.580
GMC(MALA) 0.553 0.312 0.177 0.104 0.066 0.042 0.026 0.016
MMALA 0.845 0.716 0.608 0.517 0.439 0.374 0.321 0.276
GMC(MMALA) 0.541 0.300 0.169 0.097 0.060 0.037 0.023 0.014
@@
where $c_{\gamma}=\sum_{\gamma^{\prime}\in\mathcal{N}(\gamma)}\psi(\gamma|z)$. For implementing this proposed algorithm, we need to efficiently compute $\psi(\gamma|z)$ for all $\gamma^{\prime}\in\mathcal{N}(\gamma)$. Note that Zhou et al. (2022) consider only the ‘addition’ and ‘deletion’ moves in their simulation examples, greatly reducing the computational burden. Here, we use the fast Cholesky updates of Li et al. (2023) for rapidly computing $\psi(\gamma|z)$ for all $\gamma^{\prime}\in\mathcal{N}(\gamma)$. Once $g(\gamma^{\prime}|\gamma),\gamma^{\prime}\in\mathcal{N}(\gamma)$ are computed, the inner product $\langle\sqrt{f},\sqrt{g}\rangle$ is available. Furthermore, both Yang et al. (2016) and Zhou et al. (2022) use Zellner’s (1986) $g-$prior on $\beta$ and a different prior on $\gamma$. The $g-$prior, although a popular alternative to the independent normal prior (16b), it requires all $m\times q$ sub-matrices of $W$ have full column rank for $q\leq m-1$ and the support of the prior on $\gamma$ is restricted to models of size at most $m-1$.
@@
Our simulation experiments are conducted using 100 simulated pairs of training and testing datasets. For each simulation setting, we set $p=10000$ and $m=400$ for both training and testing data sets. Following Li et al. (2023), we choose $w$ and $\lambda$ to be $\sqrt{m}/p$ and $m/p^{2}$, respectively. The error variance $\sigma^{2}$ is set by assuming different theoretical $R^{2}$ values. While the results for $R^{2}=90\%$ are provided in Table 7, the supplement contains the results for $R^{2}=60\%$ and $R^{2}=75\%$. All results are based on 100 iterations of the GMC chains and 50,000 iterations of the RW chains with all chains started at the null model.
@@
We first ran the GMC1 chain with $\lambda=m/p^{2}=5.61865e-$10 and $w=\sqrt{m}/p=2.37037e-$5 (the default choices for $\lambda$ and $w$ in the function geomc.vs in the accompanying R package geommc) for $N=100$ iterations starting from the null model. Only one variable resulted in with estimated MIP above 0.5. Indeed, $\hat{\gamma}_{\text{med}}=1_{83878775}$. The $R^{2}$ value based on fitting $\hat{\gamma}_{\text{med}}$ on the response $z$ is 20.07%. Next, we ran the GMC1 chain for 100 iterations starting at the null model, but this time following Li et al. (2023), we took $\lambda=\sqrt{m}=19.20937$ (high shrinkage) and a higher value for $w=m/p=0.00046$. This time, $\hat{\gamma}_{\text{med}}$ contained four variables with $\hat{\gamma}_{\text{med}}=(1_{83878775},2_{175541357},3_{226820323},8_{27127525})$. In addition to the estimated median model $\hat{\gamma}_{\text{med}}$, we also consider a weighted average model (WAM) $\hat{\gamma}_{\text{wam}}$. For the unique MCMC samples $\{\gamma^{(n)}\}_{n=1}^{N}$ we assign the weights $\text{wt}_{n}=\psi(\gamma^{(n)}|z)/\sum_{i=1}^{N}\psi(\gamma^{(i)}|z)$ according to the marginal posterior pmf (17). Then, the approximate marginal inclusion probability for the $j$th variable is computed as $\hat{\pi}_{j}=\sum_{n=1}^{N}\text{wt}_{n}\mathbb{I}(\gamma^{(n)}_{j}=1)$ and define the WAM as the model containing variables $j$ with $\hat{\pi}_{j}>0.5.$ Based on the 100 iterations of the GMC1 chain, $\hat{\gamma}_{\text{wam}}$ consisted of five variables with $\hat{\gamma}_{\text{wam}}=(1_{83878775},2_{79769999},2_{175541357},8_{27127525},10_{10606917})$. The $R^{2}$ values for $\hat{\gamma}_{\text{med}}$ and $\hat{\gamma}_{\text{wam}}$ are 39.10% and 43.93%, respectively. We analyzed the real dataset on the Linux server mentioned in Example 6, where it took 5.34 minutes to complete 100 iterations of GMC1 with $\lambda=\sqrt{m}$ and $w=m/p$. With these values of $(\lambda,w)$, we repeated 100 iterations of GMC1 50 times, each time with a different seed. The models $\hat{\gamma}_{\text{med}}$ and $\hat{\gamma}_{\text{wam}}$ vary in these 50 runs, suggesting that the posterior surface is highly multimodal. The range of the size of $\hat{\gamma}_{\text{med}}$ and $\hat{\gamma}_{\text{wam}}$ are (2,7) and (3, 9), respectively. Also, the ranges of $R^{2}$ values for $\hat{\gamma}_{\text{med}}$ and $\hat{\gamma}_{\text{wam}}$ are (27.29%, 46.85%) and (32.76%, 57.18%), respectively.
@@
When we ran GMC2 for 100 iterations with $\lambda=m/p^{2}$ and $w=\sqrt{m}/p$ started at the null model, both $\hat{\gamma}_{\text{med}}$ and $\hat{\gamma}_{\text{wam}}$ resulted in the empty model, although the marker with the largest MIP and the largest weighted MIP was the same SNP $1_{83878775}$ obtained by GMC1 with these choices for $(\lambda,w)$. Indeed, the largest MIP and weighted MIP were 0.47 and 0.49, respectively. Next, we ran the GMC2 chain for 100 iterations started at the null model with $\lambda=\sqrt{m}$ and $w=m/p$. Based on the 100 iterations of the GMC2 chain, $\hat{\gamma}_{\text{med}}$ consisted of eight variables with $\hat{\gamma}_{\text{med}}=(1_{83878648},2_{175541342},2_{179751269},4_{225874150},5_{21046482},8_{14610995},8_{72384070},$ $8_{115299982})$ and
$\hat{\gamma}_{\text{wam}}=(1_{83878648},2_{175541342},3_{170821024},4_{237392971},8_{72384070},9_{4338284})$. The $R^{2}$ values for $\hat{\gamma}_{\text{med}}$ and $\hat{\gamma}_{\text{wam}}$ in this case are 43.09% and 49.37%, respectively. It took 11.22 minutes to complete 100 iterations of GMC2 with $\lambda=\sqrt{m}$ and $w=m/p$. Finally, based on 50 repetitions of 100 iterations of the GMC2 chain with $\lambda=\sqrt{m}$ and $w=m/p$, each time with a different seed, we observe the ranges of the size of $\hat{\gamma}_{\text{med}}$ and $\hat{\gamma}_{\text{wam}}$ are (4,17) and (2, 8), respectively. Also, the ranges of $R^{2}$ values for $\hat{\gamma}_{\text{med}}$ and $\hat{\gamma}_{\text{wam}}$ are (40.11%, 60.10%) and (28.91%, 54.61%), respectively.